# Space Group Equivariant Crystal Diffusion

### v2

@@ 1 Introduction
Despite the fact that most existing crystal generative models ignore space groups and Wyckoff positions, they are critical for modeling real materials. Firstly, space groups and Wyckoff positions correlate strongly with materials properties; Neumann’s principle states that all crystal properties share the same invariances as the crystal itself [65]. Thus even slightly perturbing atoms out of special Wyckoff positions will reduce the crystal’s space group symmetry and can subsequently cause significant (even discontinuous) changes to its macroscopic properties [14, 11, 82, 9]. Secondly, we show empirically in Figure 1 that known materials are usually invariant to high symmetry space groups with atoms in zero-volume Wyckoff positions. An intuitive explanation for this phenomenon is that atoms of the same species in a given crystal are most stable in a specific neighboring environment, and the tendency to attain this environment naturally leads to symmetry [32]. Yet, most existing crystal generative models learn continuous distributions over all three spatial dimensions of atom positions, assigning zero probability measure to placing atoms in special Wyckoff positions. One might argue that crystals from these models can simply be relaxed into high symmetry positions with DFT or machine learned force fields. However, besides the apparent difficulty of generating atoms sufficiently close to high symmetry positions for relaxation [107, 31, 42], such a framework assigns different model probabilities to crystals that relax into the same structure, obfuscating training and evaluation.
@@ 3 Preliminaries
$d\mathbf{x}=[\mathbf{f}(\mathbf{x},t)-\gamma(t)^{2}\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x})]dt+\eta(t)d\mathbf{\bar{w}},$ (2)
@@ 5 Experiments
Datasets and evaluations.
We evaluated SGEquiDiff on two benchmark datasets: MP20 [100], containing 45,231 experimentally known crystals with up to 20 atoms per unit cell, and the more challenging MPTS52 dataset [4], containing 40,476 experimentally known crystals with up to 52 atoms per unit cell. Evaluations were conducted on 10,000 generated crystals. Following Xie et al. [100], we computed structural and compositional validity percentages using heuristics of interatomic distances and charge, respectively. We note that only $\sim$90% of crystals in MP20 pass the composition validity checker based on SMACT [16]. Of the 10,000 generated crystals, we randomly sampled 1,000 determined to be both structurally and compositionally valid. Of these 1,000 crystals, we determined how many were unique and novel with respect to the training data using pymatgen’s StructureMatcher [70] with stol=0.3, angle_tol=5, and ltol=0.2 (U.N. structures). We also followed SymmCD and calculated the fraction of 1,000 valid crystals with a unique and novel template (U.N. templates), where a template is defined as a space group and multiset of occupied Wyckoff positions. U.N. structures were used to compute (1) distribution distances between ground truth test and generated materials properties, including Wasserstein distances for atomic density $\rho$ and number of unique elements $N_{el}$ as well as Jensen-Shannon divergences (JSD) for space group $G$ and occupied Wyckoff dimensionalities $d_{\mathrm{Wyckoff}}$; (2) Central Moment Discrepancy (CMD) [106] up to 50 moments between ground truth test and generated crystal CrystalNN structural fingerprints [109]; and (3) structural and compositional diversity as measured by average pairwise $L_{2}$-distances between CrystalNN and Magpie [93] fingerprints, respectively. Average sampling times per batch of 500 crystals were measured on an NVIDIA A40 GPU. We list hyperparameters and training times for SGEquiDiff in Sec. A.8.
Table 1: MP-20 dataset. We report metrics of validity for 10,000 sampled crystals; uniqueness and novelty for 1,000 valid crystals; diversity and goodness-of-fit distributional distances for valid, unique, and novel crystals; and stability for 100 valid, unique, and novel crystals. Stability is defined as having a DFT-predicted $E_{\mathrm{hull}}$ < 0.1 eV/atom with respect to the Materials Project v2022.8.23. $T$ represents the inference-time temperature of discrete distributions in SGEquiDiff. Unless noted otherwise, TT=1.0. Best result is bolded, second best is underlined. Standard deviations over three training runs are reported in parentheses as errors of rightmost digits.
@@
Finally, for 100 random U.N. structures, we conducted structure relaxations with expensive density functional theory (DFT) calculations and assessed whether the resulting crystals were thermodynamically stable. Aligned with analyses of DFT-calculated energies for experimentally observed crystals [88] and of DFT errors relative to experiment [87, 5], stability was defined as having a predicted energy above the hull less than 0.1 eV/atom with respect to crystals in the Materials Project v2022.8.23 [39]. Further details on the DFT calculations are in Sec. A.5. The fraction of stable materials was multiplied by the fraction of U.N. structures to calculate the stable, unique, and novel (S.U.N.) rate. We qualify that a low energy above the hull as predicted by DFT has been found to be a necessary but insufficient condition for synthesizability [88]; developing physical theories for predictive synthesis is an active area of research in the materials science community [60, 25, 43, 12, 68, 48]. Due to the computational expense of DFT calculations, we only computed S.U.N. rates on a subset of baseline models. While prior works pre-relax generated crystals with foundational machine learning interatomic potentials (MLIPs) before conducting DFT relaxations [62, 86, 107], we opted not to use MLIPs to avoid conflating their biases [17, 53] with our evaluations of the generative models.
@@ Results.
On MP20, we found SGEquiDiff was competitive on all metrics and notably achieved the highest S.U.N. rate (Table 1). Unsurprisingly, we found that SGEquiDiff and other space group-constrained models (DiffCSP++, SymmCD) strongly outperformed baselines on the JSD metric for space groups and Wyckoff dimensions. We also show that by increasing the inference-time temperature of SGEquiDiff’s categorical distributions over space group, Wyckoff position, element, and the stop token, the U.N. rates and composition diversity metric can be significantly improved at the expense of slightly lower validity rates and higher distributional distances to the test set. Inspecting normalized root mean square Cartesian displacements between as-generated and DFT-relaxed structures, we found all benchmarked models performed similarly, with MatterGen being the best on average (Table 4). We note that MatterGen (44.6M parameters) is significantly larger than SGEquiDiff (5.5M parameters). Examples of S.U.N. crystals from SGEquiDiff are shown in Fig. 4.
We also benchmarked SGEquiDiff on the more difficult and memory-intensive MPTS52 dataset (Table 2), which most other models do not present results on. Using the same hyperparameters as used on MP20, we found that SGEquiDiff was able to scale to more atoms while maintaining a high rate of valid, unique, and novel crystals. SGEquiDiff notably achieved the highest U.N. template rate, composition validity, and Wasserstein distances for density and number of elements. Interestingly, all space group-constrained models significantly outperformed DiffCSP in the validity metrics, highlighting the advantage of enforcing space group symmetry as an inductive bias.
@@
To isolate effects of various components of SGEquiDiff, we performed ablations on the MP-20 dataset. First, we replaced our autoregressive lattice sampler with the DDPM-based [33] sampling of E(3)-invariant lattice matrix representations proposed in DiffCSP++ and used in SymmCD (Table 6). We used the same cosine noise scheduler as DiffCSP++, jointly diffusing lattices and atom coordinates. We term this variant of SGEquiDiff as +LDiff. We observed that +LDiff improved the Wasserstein distance for number of elements (possibly because the conditioning of the Wyckoff-Element transformer on lattices was removed) and CMD of structural fingerprints. However, our autoregressive lattice sampler yielded better structural validity, composition validity, JSD of Wyckoff dimensionalities, and S.U.N. rate. We also observed that in roughly $1\%$ of samples, the +LDiff model produced lattice matrices with negative determinant, resulting in a spurious inversion operation that does not preserve the symmetry of the 22 chiral space groups. We show examples of S.U.N. crystals from +LDiff in Fig. 5
We also replaced our space group and Wyckoff position features with those introduced in SymmCD [55] (Table 5). We found our featurizations yielded better structure validity, composition validity, Wasserstein distance for atomic density, structural divergence, and validation log-likelihoods of ground truth Wyckoff positions. In contrast, the SymmCD features yielded more diverse samples, evidenced by higher U.N. rates and composition diversity.
@@ 6 Conclusion
Due to computational constraints, we were only able to conduct 100 DFT relaxations per model, adding unknown variance to our reported S.U.N. rates. SGEquiDiff’s training data was pre-processed to assign space groups to crystals with nonzero tolerances over atom positions and lattice angles; it is possible that some materials were incorrectly symmetrized to higher symmetry space groups than representative of reality. SGEquiDiff also does not provide guidance for how to synthesize generated materials. Other symmetry groups relevant to crystals were ignored, including magnetic space groups [94], spin space groups [98], and layer groups [23]. Practically relevant deviations of real materials from perfect crystals were not considered, including defects, compositional disorder, surface effects, and interfacial effects. Future work might include application to a broader set of tasks such as those derived from non-scalar properties [102, 59, 95] or materials spectroscopy [56]. Furthermore, training non-equivariant models with space group invariant target distributions and/or applying inference-time space group symmetrization may enable a broader diversity of model architectures to generate high symmetry crystals.
@@ A.4.2 Graph Neural Network
A.5 Density Functional Theory Calculations
Total energy calculations were performed with density functional theory (DFT) using the Vienna Ab Initio Simulation Package (VASP) [54] with Projector Augmented Wave (PAW) pseudopotentials [8]. The Perdew-Burke-Ernzerhof (PBE) [73] exchange-correlation functional was used for structural relaxations. All input parameters, plane wave cutoffs, convergence criteria, k-point densities were determined by pymatgen MPRelaxSet [69]. For direct comparison with Materials Project (MP) data, we applied the MP’s energy correction scheme for anions and mixing GGA/GGA+U calculations [37, 38]. Crystal stabilities were determined through convex hull analyses using all competing phases available in MP v2022.8.23 [39].
A.5.1 Space group and Wyckoff position features
@@ A.6 Ablation results
Table 6: LDiff ablation on MP-20. We replaced our autoregressive lattice sampling method with joint diffusion of atom coordinates and E(3)-invariant lattice matrix representations proposed in DiffCSP++ [41] (+LDiff). We report metrics of validity for 10,000 sampled crystals (including the percentage of generated lattices with positive determinant, a requirement to enforce symmetry of chiral space groups); uniqueness and novelty for 1,000 valid crystals; diversity and goodness-of-fit distributional distances for valid, unique, and novel crystals; and stability for 100 valid, unique, and novel crystals. Stability is defined as having a DFT-predicted $E_{\mathrm{hull}}$ < 0.1 eV/atom with respect to the Materials Project v2022.8.23. Best result is bolded, second best is underlined. Standard deviations over three training runs are reported in parentheses as errors of rightmost digits.
@@
+LDiff 227(12) 98.95(49) 97.02(97) 84.10(63) 20.3(20) 80.3(10) 0.1435(613) 0.1292(86) 0.1847(122) 0.0579(151) 0.1549(72) 0.9474(131) 15.87(17) 13.65
A.7 Examples of S.U.N. crystals

### v3

@@ 1 Introduction
Despite the fact that most existing crystal generative models ignore space groups and Wyckoff positions, they are critical for modeling real materials. Firstly, space groups and Wyckoff positions correlate strongly with materials properties; Neumann’s principle states that all crystal properties share the same invariances as the crystal itself [65]. Thus even slightly perturbing atoms out of special Wyckoff positions will reduce the crystal’s space group symmetry and can subsequently cause significant (even discontinuous) changes to its macroscopic properties [14, 11, 82, 9]. Secondly, we show empirically in Figure 1 that known materials are usually invariant to high symmetry space groups with atoms in zero-volume Wyckoff positions. An intuitive explanation for this phenomenon is that atoms of the same species in a given crystal are most stable in a specific neighboring environment, and the tendency to attain this environment naturally leads to symmetry [32]. Yet, most existing crystal generative models learn continuous distributions over all three spatial dimensions of atom positions, assigning zero probability measure to placing atoms in special Wyckoff positions. One might argue that crystals from these models can be relaxed into high symmetry positions with DFT or machine learned force fields. However, besides the apparent difficulty of generating atoms sufficiently close to high symmetry positions for relaxation [107, 31, 42], such a framework assigns different model probabilities to crystals that relax into the same structure, obfuscating training and evaluation.
@@ 3 Preliminaries
$d\mathbf{x}=[\mathbf{f}(\mathbf{x},t)-\eta(t)^{2}\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x})]dt+\eta(t)d\mathbf{\bar{w}},$ (2)
@@ 5 Experiments
Evaluations.
We evaluated SGEquiDiff on two benchmark datasets: MP20 [100], containing 45,231 experimentally known crystals with up to 20 atoms per unit cell, and the more challenging MPTS52 dataset [4], containing 40,476 experimentally known crystals with up to 52 atoms per unit cell. Evaluations were conducted on 10,000 generated crystals. Following Xie et al. [100], we computed structural and compositional validity percentages using heuristics of interatomic distances and charge, respectively. We note that only $\sim$90% of crystals in MP20 pass the composition validity checker based on SMACT [16]. Of the 10,000 generated crystals, we randomly sampled 1,000 determined to be both structurally and compositionally valid. Of these 1,000 crystals, we determined how many were unique and novel with respect to the training data using pymatgen’s StructureMatcher [70] with stol=0.3, angle_tol=5, and ltol=0.2 (U.N. structures). We also followed SymmCD and calculated the fraction of unique and novel templates (U.N. templates), where a template is defined as a space group and multiset of occupied Wyckoff positions. U.N. structures were used to compute (1) distribution distances between ground truth test and generated materials properties, including Wasserstein distances for atomic density $\rho$ and number of unique elements $N_{el}$ as well as Jensen-Shannon divergences (JSD) for space group $G$ and occupied Wyckoff dimensionalities $d_{\mathrm{Wyckoff}}$; (2) Central Moment Discrepancy (CMD) [106] up to 50 moments between ground truth test and generated CrystalNN structural fingerprints [109]; and (3) structural and compositional diversity as measured by average pairwise $L_{2}$-distances between CrystalNN and Magpie [93] fingerprints, respectively. Average sampling times per batch of 500 crystals were measured on an NVIDIA A40 GPU. We list hyperparameters and training times for SGEquiDiff in Sec. A.7.
Table 1: MP-20 dataset. We report metrics of validity for 10,000 sampled crystals; uniqueness and novelty for 1,000 valid crystals; diversity and goodness-of-fit distributional distances for valid, unique, and novel crystals; and stability for valid, unique, and novel crystals. Stability is defined as having a DFT-predicted $E_{\mathrm{hull}}$ < 0.1 eV/atom with respect to the Materials Project v2022.8.23. $T$ represents the inference-time temperature of discrete distributions in SGEquiDiff. Unless noted otherwise, TT=1.0. Best result is bolded, second best is underlined. Standard deviations over three training runs are reported in parentheses as errors of rightmost digits.
@@
Finally, for all U.N. structures out of 1,000 random valid crystals, we conducted structure relaxations with expensive density functional theory (DFT) calculations to assess thermodynamic stability. Aligned with analyses of DFT-calculated energies for experimentally observed crystals [88] and of DFT errors relative to experiment [87, 5], stability was defined as having a predicted energy above the hull less than 0.1 eV/atom with respect to crystals in the Materials Project v2022.8.23 [39]. Further details on the DFT calculations are in Sec. A.8. The fraction of stable materials was multiplied by the fraction of U.N. structures to calculate the stable, unique, and novel (S.U.N.) rate. We qualify that a low energy above the hull as predicted by DFT has been found to be a necessary but insufficient condition for synthesizability [88]; developing physical theories for predictive synthesis is an active area of research in the materials science community [60, 25, 43, 12, 68, 48]. Due to the computational expense of DFT calculations, we only computed S.U.N. rates on a subset of baseline models. While prior works pre-relax generated crystals with foundational machine learning interatomic potentials (MLIPs) before conducting DFT relaxations [62, 86, 107], we opted not to use MLIPs to avoid conflating their biases [17, 53] with our evaluations of the generative models.
@@ Results.
On MP20, we found SGEquiDiff was competitive on all metrics and achieved the highest S.U.N. rate (Table 1). While MatterGen achieved a comparable S.U.N. rate, we note that it is a significantly larger model (44.6M parameters) than SGEquiDiff (5.5M parameters), sampled crystals $\sim$16 times slower, and yielded worse JSD metrics for space groups and Wyckoff dimensions. Compared to DiffCSP++, SGEquiDiff achieved a slightly higher S.U.N. rate and significantly higher U.N. template rate with $\sim$2.2 times fewer model parameters. Unsurprisingly, SGEquiDiff and other space group-constrained models (DiffCSP++, SymmCD) strongly outperformed baselines on the JSD metric for space groups and Wyckoff dimensions. We also show that by increasing the inference-time temperature of SGEquiDiff’s categorical distributions over space group, Wyckoff position, element, and the stop token, the U.N. rates and composition diversity metric can be significantly improved at the expense of slightly lower validity rates and higher distributional distances to the test set. Inspecting normalized root mean square Cartesian displacements between as-generated and DFT-relaxed structures, we found all models performed similarly, with MatterGen being the best on average (Table 4). Examples of S.U.N. crystals from SGEquiDiff are shown in Fig. 4.
We also benchmarked SGEquiDiff on the more difficult MPTS52 dataset (Table 2), which most other models do not present results on. Using the same hyperparameters as used on MP20, we found that SGEquiDiff was able to scale to more atoms while maintaining a high rate of valid, unique, and novel crystals. SGEquiDiff notably achieved the highest U.N. template rate, composition validity, and Wasserstein distances for density and number of elements. Interestingly, all space group-constrained models significantly outperformed DiffCSP in the validity metrics, highlighting the advantage of enforcing nontrivial space group symmetries as an inductive bias.
@@
To isolate effects of various components of SGEquiDiff, we performed ablations on the MP-20 dataset. First, we replaced our autoregressive lattice sampler with the DDPM-based [33] sampling of E(3)-invariant lattice matrix representations proposed in DiffCSP++ and used in SymmCD (Table 6). We used the same cosine noise scheduler as DiffCSP++, jointly diffusing lattices and atom coordinates. We term this variant of SGEquiDiff as +LDiff. We observed that +LDiff improved the Wasserstein distance for number of elements (possibly because conditioning of the Wyckoff-Element transformer on lattices was removed) and CMD of structural fingerprints. However, our autoregressive lattice sampler yielded better validity, JSD of Wyckoff dimensionalities, and S.U.N. rate. We also observed that in roughly $1\%$ of samples, the +LDiff model produced lattice matrices with negative determinant, resulting in a spurious inversion that does not preserve the symmetry of the 22 chiral space groups. We show examples of S.U.N. crystals from +LDiff in Fig. 5
We also replaced our space group and Wyckoff position features with those introduced in SymmCD [55] (Table 5). We found our featurizations yielded better structure validity, composition validity, Wasserstein distance for atomic density, structural divergence, and validation log-likelihoods of ground truth Wyckoff positions. In contrast, SymmCD features yielded more diverse samples, evidenced by higher U.N. rates and composition diversity.
@@ 6 Conclusion
Due to computational constraints, we were only able to calculate S.U.N. rates on 1,000 valid crystals per model, adding unknown variance to reported metrics. SGEquiDiff’s training data was pre-processed to assign space groups to crystals with nonzero tolerances over atom positions and lattice angles; it is possible that some materials were incorrectly symmetrized to higher symmetry space groups than representative of reality. SGEquiDiff also does not provide guidance for how to synthesize generated materials. Other symmetry groups relevant to crystals were ignored, including magnetic space groups [94], spin space groups [98], and layer groups [23]. Practically relevant deviations of real materials from perfect crystals were not considered, including defects, compositional disorder, surface effects, and interfacial effects. Future work might include application to a broader set of tasks such as those derived from non-scalar properties [102, 59, 95] or materials spectroscopy [56]. Furthermore, training non-equivariant models with space group invariant target distributions and/or applying inference-time space group symmetrization may enable a broader diversity of model architectures to generate high symmetry crystals.
@@ A.4.2 Graph Neural Network
A.4.3 Space group and Wyckoff position features
@@ A.5 Ablation results
Table 6: LDiff ablation on MP-20. We replaced our autoregressive lattice sampling method with joint diffusion of atom coordinates and E(3)-invariant lattice matrix representations proposed in DiffCSP++ [41] (+LDiff). We report metrics of validity for 10,000 sampled crystals (including the percentage of generated lattices with positive determinant, a requirement to enforce symmetry of chiral space groups); uniqueness and novelty for 1,000 valid crystals; diversity and goodness-of-fit distributional distances for valid, unique, and novel crystals; and stability for valid, unique, and novel crystals. Stability is defined as having a DFT-predicted $E_{\mathrm{hull}}$ < 0.1 eV/atom with respect to the Materials Project v2022.8.23. Best result is bolded, second best is underlined. Standard deviations over three training runs are reported in parentheses as errors of rightmost digits.
@@
+LDiff 227(12) 98.95(49) 97.02(97) 84.10(63) 20.3(20) 80.3(10) 0.1435(613) 0.1292(86) 0.1847(122) 0.0579(151) 0.1549(72) 0.9474(131) 15.87(17) 13.65*
* For efficiency, stability calculations were conducted on 100 random valid, unique, and novel structures.
A.6 Examples of S.U.N. crystals